# Straightening the Ruler: Field-Level Inference of the BAO Scale with LEFTfield

### v1

@@ 3 Code Implementation
In this section we briefly summarize the way the theory is implemented in the LEFTfield code. We start from the sampling of initial conditions. We place those on a grid of size $N_{g}^{\Lambda}$, and multiply them with the square root of the linear power spectrum to obtain the initial field. The size of the initial grid, $N_{g}^{\Lambda}$, is determined by the requirement to represent all Fourier modes (for the given box size $L$) up to the cutoff $\Lambda$. Next, we find the bias operators. This procedure slightly varies between the Eulerian and Lagrangian case.
In the Eulerian case, we construct the LPT displacement to second order (2LPT). During the construction, to make sure all physical modes are accounted for, we employ a grid of size $N_{g}^{n\Lambda}$, where $n=2$ is the maximum perturbative order. We then displace a uniform grid of pseudoparticles to Eulerian space. These pseudoparticles are then assigned (see below) to obtain the Eulerian density field $\delta$. This field is subsequently sharp-$k$ filtered as described above, after which the Eulerian bias operators are constructed. We resize the grid appropriately to ensure that all modes entering the likelihood are safe from aliasing.
In the Lagrangian bias case, the set of operators listed in Eq. (2.16) is constructed concurrently with the 2LPT displacement. In addition to the uniform-weight grid used to obtain the Eulerian matter density, we also assign particles with weights given by the $O_{L}(\mathbf{q})$, effectively displacing the Lagrangian operators to Eulerian space [22].
The assignment schemes implemented in LEFTfield are nearest-grid-point (NGP), cloud-in-cell (CIC), triangular-shaped cloud (TSC) and non-uniform-to-uniform discrete Fourier transform (NUFFT) [41], which we employ as an assignment scheme. For the forward model in this paper, we use the NUFFT, as it converges rapidly with Eulerian grid size thanks to the kernel deconvolution. NUFFT implements the non-uniform-to-uniform discrete Fourier transform [$f(\mathbf{x})\to\tilde{f}(\mathbf{k})$] by assigning particle positions $\mathbf{x}_{i}$ with weights $f(\mathbf{x}_{i})$ to a supersampled grid (increasing the resolution by a factor typically in the range 1.2 to 2) using a suitable assignment kernel with compact support (of roughly 4 to 16 grid cells). It then performs an FFT on the supersampled grid, deconvolves the assignment kernel, and finally performs a grid reduction in Fourier space to yield the desired uniform discrete Fourier transform $\tilde{f}(\mathbf{k})$. This method is approximate, but accuracy close to machine precision can be obtained for very reasonable computational effort. For the Eulerian grid size, we choose $N_{g}^{\rm Eul}$ such that the Nyquist frequency corresponds to $(3/2)\Lambda$, at which point the density assignment is converged to next-to-leading order [42].
@@ 4 Synthetic Data sets
The data sets used in this paper were generated using the LEFTfield code, which we denote as Mock A and Mock B. Both data sets were generated in a box of side length $L=2000\,h^{-1}$pc and with fiducial cosmology fixed to $\sigma_{8}=0.85$, $\Omega_{m}=0.3$, $\Omega_{\Lambda}=0.7$, $h=0.7$ and $n_{s}=0.967$. We emphasize that both mocks contain nonlinear information, as they were generated using second-order LPT, while the distinguishing factor between them lies in the choice of the bias model. Mock A was created using the second-order Lagrangian bias, while Mock B utilized the second-order Eulerian bias. For the bias parameters, we adopted values obtained through fixed-phase inference on realizations of halo catalogues (see for example [43]), specifically the halo mass range $\log_{10}(M/h^{-1}M_{\odot})=13.0-13.5$ at redshift $z=0$. In order to increase the signal to noise ratio, we used a slightly lower noise level than in the halos sample for both mocks. Ground truth $\beta_{0}$ values were deliberately selected to be different from 1 to verify unbiased inference even in the case of a nontrivial rescaling of the fiducial power spectrum. Furthermore, the mocks were generated at the cutoff $\Lambda_{0}=0.3\,h\,\text{Mpc}^{-1}$, while the inference was performed at lower $\Lambda$ values, allowing us to test the impact of model mismatch in this dimension as well. The values of all parameters used to generate these mocks are summarized in Table 1.
Mock Bias basis $P_{\epsilon}\,[\,h^{-3}\,{\rm Mpc}^{3}]$ $\beta$ $b_{\delta}$ $b_{\nabla^{2}\delta}$ $b_{\sigma\sigma}$ $b_{\text{Tr}[M^{(1)}M^{(1)}]}$ $b_{\delta^{2}}$ $b_{K^{2}}$
A Lagrangian 553.83 0.99 1.21 0 -0.26 -0.18 - -
B Eulerian 915.53 0.98 1.67 -0.07 - - -0.19 0
Table 1: Values of parameters used to produce mock datasets A and B. For both mocks, we use second order LPT combined with second order bias expansion and cutoff $\Lambda_{0}=0.3\,h\,{\rm Mpc}^{-1}$.
5 Field level BAO inference
@@ 5.1 Sampling method and data analysis
In this section, we provide a detailed account of how we sample from the posterior defined in Section 2. To achieve this, we employ a combination of the Hamiltonian Monte Carlo (HMC) sampler for initial conditions and the slice sampler for all other parameters.
@@
$\hat{s}(\textbf{x})=\begin{cases}\delta_{D}(\hat{s}-\hat{s}_{\mathrm{true}}),&\text{FixedIC},\\ \mathcal{N}(0,1),&\text{FreeIC}.\end{cases}$ (5.1)
In the case when $\hat{s}$ is fixed to ground truth, we are talking about FixedIC case. Alternatively, the case when the Gaussian prior for $\hat{s}$ is chosen, we denote as FreeIC. FixedIC case has been extensively studied in our previous work [16], while here we are focused on the FreeIC case and use the FixedIC case only as a consistency check.
The FreeIC case is challenging in particular due to the dimensionality of the parameter space of $\hat{s}$ which is of the order of $10^{6}$. To perform the sampling of such a large parameter space, we use the HMC sampler which is well suited for sampling in large dimensions. While the execution time in standard MCMC techniques typically scales linearly with the dimension of the problem $N_{\text{dim}}$, in the case of the HMC it scales as $N^{1/4}_{\text{dim}}$ [27]. Because of the way HMC generates new proposals, it requires a forward model which is differentiable with respect to the initial conditions. Fortunately, our forward model is differentiable, and the structure of the LEFTfield code allows us to find its analytical derivative by performing successive applications of the chain rule.
@@
When employing the likelihood marginalized over bias parameters, one MCMC chain used Gaussian priors for bias parameters. More specifically, the priors used were $\mathcal{N}(0.01,5)$ for $b_{\delta}$ and $\mathcal{N}(0,5)$ for the other bias parameters. In addition to this, we operated two other chains that employed uniform priors: $\mathcal{U}(0.01,10)$ for $b_{\delta}$, and $\mathcal{U}(-30,30)$ for the remaining bias parameters. However, the bias coefficients are well constrained by the data, and the posteriors are likelihood-dominated for both chains. Hence, despite the different priors, all chains are found to yield consistent posteriors in $\beta,\sigma_{\epsilon}$ and $\sigma_{\epsilon,2}$. Therefore, we aggregated the results from all three chains in the analysis. For comparison, we have also performed inferences using a likelihood that was not marginalized over bias parameters; the results of this inference are shown in App. A, and provide evidence for the above statement.
@@ 5.2 Field-level results
Figure 1: Trace plot of parameter $\beta$ for $\Lambda=0.2\,h\,{\rm Mpc}^{-1}$ for three independent MCMC chains, where $n$ denotes the sample index. The chain shown in green started from the true initial conditions $\hat{s}_{\rm true}$ while the other two chains started from random initial conditions $\hat{s}$. Each chain started from a different initial $\beta$ value, with all quickly converging to the same value. Dashed black line indicates the ground truth value $\beta_{0}$.
@@
Figure 2: The normalized auto-correlation function for parameter $\beta$ inferred at $\Lambda=0.2\,h\,{\rm Mpc}^{-1}$. In the figure we also display the correlation length value $\tau$ together with the maximum separation $M$ between the samples considered, both of which are defined in App. A. The correlation length of $\beta$ is estimated to be $\tau\simeq 244$ samples. In the lower panel, we zoom into the first 1800 samples of the chain.
@@
Our main findings for Mock A are summarized in Fig. 3. In the left panel, we show the remaining systematic bias in the $\beta$ value. For the lower cutoffs, $\Lambda=0.15\,h\,{\rm Mpc}^{-1}\text{ and }\Lambda=0.18\,h\,{\rm Mpc}^{-1}$, the bias is below 1%, however in the case of $\Lambda=0.2\,h\,{\rm Mpc}^{-1}$ the systematic offset is around 1.8%. This is most likely due to our choice to employ a second-order bias expansion for the inference. We generated the mocks at a higher cutoff, and thus they require more bias terms to describe at the lower cutoff used for the inference.
@@
The right panel of Fig. 3 displays the 68% confidence-level error bars for $\beta$, denoted as $\sigma_{\rm F}(\beta)$, obtained through the EFT likelihood across varying $\Lambda$ values. For the lowest cutoff, $\sigma_{\rm F}(\beta)$ slightly exceeds $1\%$, whereas for the two higher cutoffs, it reduces to $0.71\%$ and $0.66\%$, respectively. The reduction in error bar size with increasing cutoff was anticipated, as a larger $\Lambda$ permits the inclusion of more modes in both the forward model and the likelihood, thereby providing additional information. We can also compare the behavior of $\sigma_{\rm F}(\beta)$ as a function of $\Lambda$ with the ideal expectation from mode counting, $\sigma_{\rm F}(\beta)\propto N_{\text{mode}}^{-1/2}(\Lambda)$, finding very rough agreement. More precisely, the error bar $\sigma_{\rm F}(\beta)$ improves more rapidly than the mode scaling between $\Lambda=0.15\,h\,{\rm Mpc}^{-1}$ and $\Lambda=0.18\,h\,{\rm Mpc}^{-1}$, while between $\Lambda=0.18\,h\,{\rm Mpc}^{-1}$ and $\Lambda=0.2\,h\,{\rm Mpc}^{-1}$, it shrinks at a lower rate than estimated from mode counting. Note however that the simple mode counting ignores the varying sensitivity to a change in $\beta$ of the linear power spectrum as a function of $k$.
@@
Figure 3: Inference results for the Mock A. Left panel shows inferred BAO scale relative to ground truth obtained using Lagranigan bias for sampling. FreeIC are represented using circle marker and FixedIC using a square. On the right we show the 68% CL error bar on the BAO scale, $\sigma_{\rm F}(\beta)$, as a function of cutoff $\Lambda$.
@@
Figure 4: Inference results for Mock B. The left panel displays the inferred BAO scale relative to the ground truth, where the BAO scale is sampled alongside the initial conditions. The Lagrangian bias points, computed for the same values of $\Lambda$, have been slightly displaced horizontally for better visibility. On the right, we show the 68% CL error bar, $\sigma_{\rm F}(\beta)$, as a function of the cutoff $\Lambda$. The results obtained using the Lagrangian bias are depicted in blue, while those obtained using Eulerian bias are shown in orange.
@@
Turning our attention to Mock B, this data set is generated using Eulerian bias and sampled with both Eulerian and Lagrangian bias. The left panel of Fig. 4 shows the inferred $\beta$ for both Eulerian and Lagrangian bias cases. In the instance of Eulerian bias, we observe that the residual systematic bias remains below 1% across all considered cutoffs. Additionally, the ground truth value of $\beta$ is consistently recovered within $1\sigma$. Conversely, when using the Lagrangian bias model for inference, a slightly higher systematic bias is evident. Specifically, for $\Lambda=0.15\,h\,{\rm Mpc}^{-1}$ and $0.18\,h\,{\rm Mpc}^{-1}$, this bias hovers around 1%, while for $\Lambda=0.2\,h\,{\rm Mpc}^{-1}$, it increases somewhat to 1.7%. This is consistent with what we found for Mock A, and would likely be solved by going to a higher order in bias in the inference.
In the right panel of Fig. 4, we show the inferred 68% CL error bar $\sigma_{\rm F}(\beta)$ for the Mock B at the different cutoffs. We find that in the case of both bias models the size of the error bar is shrinking with increasing $\Lambda$. This trend aligns with our expectations, as a larger $\Lambda$ implies the inclusion of more modes in both the forward model and the likelihood.
Comparing the error bar size, $\sigma_{\rm F}(\beta)$, between the Eulerian and Lagrangian models shows that the error bar is smaller in the case of the Lagrangian bias model. This difference is likely attributable to the construction of bias operators in these models. In the Eulerian bias model, bias operators are constructed using the filtered matter field, $\delta_{\Lambda}$. This second filter (in addition to the one applied to the initial conditions) removes some mode-coupling contributions that are under control and kept in the Lagrangian bias model, which only filters the initial conditions (both bias models of course use the same likelihood filter). The downside of the Lagrangian approach is the higher computational cost due to the additional density assignments needed.
In Fig. 5, we show posteriors of parameters sampled on the Mock B dataset using $\Lambda=0.18\,h\,{\rm Mpc}^{-1}$. The two panels correspond to the Eulerian and Lagrangian biases, respectively. Mock B was generated at $\Lambda_{0}=0.3\,h\,{\rm Mpc}^{-1}$ with $\sigma_{\epsilon}=0.9$. From Eq. (2.19), we find that this will correspond $\sigma_{\epsilon}\approx 0.41$ at $\Lambda=0.18\,h\,{\rm Mpc}^{-1}$. We see that the inferred value of $\sigma_{\epsilon}$ is indeed consistent with the expected value within errors. However, we notice that $\sigma_{\epsilon,2}$ is now larger than zero. This elevation in noise level is due to the fact that inference is performed at $\Lambda<\Lambda_{0}$. Since the model is nonlinear, the modes between $\Lambda$ and $\Lambda_{0}$ which are integrated out lead to additional effective noise contributions, as shown explicitly in [23] and at the level of renormalization-group equations in [45].
@@
Figure 5: Posterior of parameters for the FreeIC inference in the case of Mock B. Left panel (a) represents the Eulerian bias model while the right panel (b) represents the Lagrangian bias. The dotted gray line indicates the ground truth value $\beta_{0}$ in each case. The inference was performed at $\Lambda=0.18\,h\,{\rm Mpc}^{-1}$ in both cases.
6 Comparing the field-level results to the reconstruction approach

### v2

@@ 3 Code Implementation
In this section we briefly summarize how the forward model and likelihood are implemented in the LEFTfield code. We start from the sampling of initial conditions. The initial conditions, parametrized via the field $\hat{s}$, are discretized on a grid of size $N_{g}^{\Lambda}$, and the linear density field is obtained via Eq. (2.2). The size of the initial grid, $N_{g}^{\Lambda}$, is determined by the requirement to represent all Fourier modes (for the given box size $L$) up to the cutoff $\Lambda$. Next, we construct the bias operators. This procedure slightly varies between the Eulerian and Lagrangian case.
In the Eulerian case, we construct the LPT displacement to second order (2LPT). During the construction, all physical modes are accounted for by suitably enlarging the grid size (see [42] for a detailed discussion). We then displace a uniform grid of pseudoparticles to Eulerian space. These pseudoparticles are then assigned (see below) to obtain the Eulerian density field $\delta$. This field is subsequently sharp-$k$ filtered as described above, after which the Eulerian bias operators are constructed. We resize the grid appropriately to ensure that all modes entering the likelihood are safe from aliasing.
In the Lagrangian bias case, the set of operators listed in Eq. (2.16) is constructed concurrently with the 2LPT displacement. In addition to the uniform-weight grid used to obtain the Eulerian matter density, we also assign particles with weights given by the $O_{L}(\bm{q})$, effectively displacing the Lagrangian operators to Eulerian space [23].
The assignment schemes implemented in LEFTfield are nearest-grid-point (NGP), cloud-in-cell (CIC), triangular-shaped cloud (TSC) and non-uniform-to-uniform discrete Fourier transform (NUFFT) [43], which we employ as an assignment scheme. For the forward model in this paper, we use the NUFFT, as it converges rapidly with Eulerian grid size thanks to the kernel deconvolution. NUFFT implements the non-uniform-to-uniform discrete Fourier transform [$f(\bm{x})\to\tilde{f}(\bm{k})$] by assigning particle positions $\bm{x}_{i}$ with weights $f(\bm{x}_{i})$ to a supersampled grid (increasing the resolution by a factor typically in the range 1.2 to 2) using a suitable assignment kernel with compact support (of roughly 4 to 16 grid cells). It then performs an FFT on the supersampled grid, deconvolves the assignment kernel, and finally performs a grid reduction in Fourier space to yield the desired uniform discrete Fourier transform $\tilde{f}(\bm{k})$. This method is approximate, but accuracy close to machine precision can be obtained for very reasonable computational effort. For the Eulerian grid size, we choose $N_{g}^{\rm Eul}$ such that the Nyquist frequency corresponds to $(3/2)\Lambda$, at which point the density assignment is converged to next-to-leading order [42].
@@ 4 Synthetic Data sets
The data sets used in this paper were generated using LEFTfield, and denoted as Mock A and Mock B. Both data sets were generated in a box of side length $L=2000\,h^{-1}$pc and with a fiducial Euclidean $\Lambda$CDM cosmology characterized by $\sigma_{8}=0.85$, $\Omega_{m}=0.3$, $\Omega_{\Lambda}=0.7$, $h=0.7$ and $n_{s}=0.967$. We emphasize that both mocks contain nonlinear information, as they were generated using second-order LPT, while the distinguishing factor between them lies in the choice of the bias model. Mock A was created using the second-order Lagrangian bias, while Mock B utilized the second-order Eulerian bias. For the bias parameters, we adopted values obtained through fixed-phase inference on realizations of halo catalogues (see for example [44]), specifically the halo mass range $\log_{10}(M/h^{-1}M_{\odot})=13.0-13.5$ at redshift $z=0$. The ground-truth values for the BAO scale, $\beta_{0}$, were deliberately selected to be different from 1 to verify unbiased inference even in the case of a nontrivial rescaling of the fiducial power spectrum. Furthermore, the mocks were generated at the cutoff $\Lambda_{0}=0.3\,h\,\text{Mpc}^{-1}$, while the inference was performed at lower $\Lambda$ values, allowing us to test the impact of model mismatch in this dimension as well. The values of all parameters used to generate these mocks are summarized in Table 1.
We determine the value of $\sigma_{\epsilon}$ in the generation of the mock data by requiring that the inferred noise level, at the substantially lower cutoff scales used in the inference, is roughly comparable to that of the halo sample. The inferred noise power spectrum is listed in Table 1. The reason for this procedure is that the noise amplitude runs very strongly with cutoff [45]. The actual noise power spectra employed at the scale $\Lambda=0.3\,h\,{\rm Mpc}^{-1}$ in the mock generation is at the level of $[600-900]\,h^{-3}\,{\rm Mpc}^{3}$.
Mock Bias basis $\beta$ $b_{\delta}$ $b_{\nabla^{2}\delta}$ $b_{\sigma\sigma}$ $b_{\text{Tr}[M^{(1)}M^{(1)}]}$ $b_{\delta^{2}}$ $b_{K^{2}}$ $P_{\epsilon}^{\Lambda=0.18\,h\,{\rm Mpc}^{-1}}$
A Lagrangian 0.99 1.21 0 -0.26 -0.18 - - 1510.1
B Eulerian 0.98 1.67 -0.07 - - -0.19 0 924.44
Table 1: Values of parameters used to produce mock datasets A and B, with the exception of $P_{\epsilon}$ for which we list the inferred value at $\Lambda=0.18\,h\,{\rm Mpc}^{-1}$ in units of $\,h^{-3}\,{\rm Mpc}^{3}$ (see text for a discussion). For both mocks, we use second order LPT combined with second order bias expansion and cutoff $\Lambda_{0}=0.3\,h\,{\rm Mpc}^{-1}$.
5 Field-level BAO inference
@@ 5.1 Sampling method and data analysis
In this section, we provide a detailed account of how we sample from the posterior defined in Section 2. We employ a combination of the Hamiltonian Monte Carlo (HMC) sampler for initial conditions, coupled with the slice sampler for the BAO scale and noise parameters in a block sampling algorithm. The bias parameters on the other hand are analytically marginalized over [19].
@@
$\hat{s}(\textbf{x})=\begin{cases}\delta_{D}^{(N_{g}^{3})}(\hat{s}-\hat{s}_{\mathrm{true}}),&\text{FixedIC},\\ \mathcal{N}(0,1),&\text{FreeIC}\,,\end{cases}$ (5.1)
where $\hat{s}_{\rm true}$ denotes the ground-truth initial conditions, i.e. those used in the mock generation. The FixedIC case has been extensively studied in our previous work [16], while here we are focused on the FreeIC case and use the FixedIC case only as a consistency check.
The FreeIC case is challenging in particular due to the dimensionality of the parameter space of $\hat{s}$, which is of the order of $10^{6}$. To perform the sampling of such a large parameter space, we use the HMC sampler which is well suited for sampling in large dimensions. The computational effort required to generate a given number of independent samples scales roughly exponentially with the dimension of the problem $N_{\text{dim}}$ in the case of classical Metropolis-Hastings MCMC; in the case of well-tuned HMC sampling, it only scales as $N^{1/4}_{\text{dim}}$ [29]. Because of the way HMC generates new proposals, it requires a forward model which is differentiable with respect to the initial conditions. Fortunately, our forward model is differentiable, and the structure of the LEFTfield code allows us to find its analytical derivative by performing successive applications of the chain rule.
@@
When employing the likelihood marginalized over bias parameters, one MCMC chain used Gaussian priors for bias parameters. More specifically, the priors used were $\mathcal{N}(1,5)$ for $b_{\delta}$ and $\mathcal{N}(0,5)$ for the other bias parameters. In addition to this, we ran two other chains that employed uniform priors: $\mathcal{U}(0.01,10)$ for $b_{\delta}$, and $\mathcal{U}(-30,30)$ for the remaining bias parameters. We find that the bias coefficients are well constrained by the data, and the posteriors are likelihood-dominated for both chains. Hence, despite the different priors, all chains are found to yield consistent posteriors in $\beta,\sigma_{\epsilon}$ and $\sigma_{\epsilon,2}$. Therefore, we aggregated the results from all three chains in the analysis. For comparison, we have also performed inferences that explicitly sampled the bias parameters. The results are shown in Figs. 14–fig: Mock A Fourier in App. A, and provide evidence for the above statement.
@@
Mock Forward model $P_{\epsilon}\,[\,h^{-3}\,{\rm Mpc}^{3}]$ $\beta$ $b_{\delta}$
C Linear 271.12 0.98 1.00
D $1^{\rm st}$ order LPT 271.12 0.98 1.22
Table 2: Parameter value used to produce mock datasets C and D for validation tests. Both datasets assume a cutoff $\Lambda_{0}=0.18\,h\,{\rm Mpc}^{-1}$.
5.2 Validation tests
To validate the reliability of both our field-level results and power spectrum-based results, we conducted a series of tests by performing both inferences on simple mock data with well-defined expected outcomes. The details and results of these tests are presented briefly here. Both tests are performed without any model mismatch, and cutoffs $\Lambda=\Lambda_{0}=0.18\,h\,{\rm Mpc}^{-1}$.
5.2.1 Linear model
We first generated a mock from a linear forward model, where the initial density field is linearly extrapolated to $z=0$, and multiplied with a single linear bias coefficient $b_{\delta}$. Since the field remains linear throughout its evolution, all information about the field at $z=0$ is fully captured by the power spectrum. This implies that inference based solely on the power spectrum, combined with the fitting and MCMC sampling which we describe in Sec. 6.1, should yield the same uncertainty in $\beta$ as the field-level approach. In other words, the error bars $\sigma(\beta)$ obtained from both inference methods should be identical. We refer to this mock as Mock C, and its details are listed in Table 2.
As can be seen from Table 3, both methods give the exact same error bar, validating that the field-level analysis chain does not underestimate the error bar.
Mean and 68% CL error on $\beta$
Field-level inference Power spectrum inference
Mock C (linear) 0.9760 $\pm$ 0.0060 0.9812 $\pm$ 0.0060 (pre-rec.)
Mock D (1LPT) 0.9766 $\pm$ 0.0065 0.9773 $\pm$ 0.0066 (post-rec.)
Table 3: Results of validation tests performed on Mocks C (linear) and D (1LPT). In each case, the error bars for field-level inference agrees with the power spectrum (Mock C) and post-reconstruction power spectrum (Mock D), respectively.
5.2.2 1LPT model
The BAO reconstruction technique relies on the Zel’dovich approximation and the linear bias model to reverse the non-linear damping of the BAO feature. Consequently, if we analyze a data set generated using the Zel’dovich approximation and a linear bias model, the post-reconstruction power spectrum should recover the same information as the field-level analysis. To test this, we constructed a mock catalog, Mock D, using first-order Lagrangian perturbation theory combined with the linear bias model. The coefficients for this mock catalog are presented in Table 2 as well. This comparison provides a nontrivial test of both the field-level inference analysis chain as well as our implementation of BAO reconstruction. The latter, described in detail in Sec. 6, involves some subtleties owing to how our mock data are generated.
The results of this test are presented in Table 3. We see that the error bars in both cases are almost identical, indicating that the reconstruction procedure is both effective and consistent with the field-level inference, as expected in this case.
@@ 5.3 Field-level results
Figure 1: Trace plot of the BAO scale parameter $\beta$ for Mock A and $\Lambda=0.2\,h\,{\rm Mpc}^{-1}$ for three independent MCMC chains, where $n$ denotes the sample index. The chain shown in green started from the true initial conditions $\hat{s}_{\rm true}$ while the other two chains started from random initial conditions. Each chain started from a different initial $\beta$ value, with all quickly converging to the same value. Dashed black line indicates the ground truth value $\beta_{0}$.
@@
Figure 2: The normalized auto-correlation function for the BAO scale parameter $\beta$ inferred from Mock A at $\Lambda=0.2\,h\,{\rm Mpc}^{-1}$ (upper panel). We also display the correlation length value $\tau$ together with the maximum separation $M$ between the samples considered, both of which are defined in App. A. The correlation length of $\beta$ is estimated to be $\tau\simeq 250$ samples. In the lower panel, we zoom into the first 1800 samples of the chain.
@@
Our main findings for Mock A are summarized in Fig. 3. In the left panel, we show the inferred BAO scale, relative to the ground truth $\beta_{0}$. For the lower cutoffs, $\Lambda=0.15\,h\,{\rm Mpc}^{-1}$ and $\Lambda=0.18\,h\,{\rm Mpc}^{-1}$, the bias is below 1%, however in the case of $\Lambda=0.2\,h\,{\rm Mpc}^{-1}$ the systematic offset is around 1.8%. This is most likely due to our choice of a second-order bias expansion for the inference. The mocks were generated at a higher cutoff, which means that higher-order bias terms are effectively generated at the lower cutoff used in the inference. Moreover, higher-order bias terms are expected to become more important for higher values of $\Lambda$. To test this assumption, we repeated the inference procedure for this mock and $\Lambda=0.2\,h\,{\rm Mpc}^{-1}$ with a third-order Lagrangian bias expansion. In this case, the remaining systematic bias in $\beta$ is reduced to less than 1%.
@@
The right panel of Fig. 3 displays the 68% confidence-level error bars for $\beta$, denoted as $\sigma_{\rm F}(\beta)$, obtained through the EFT likelihood across varying $\Lambda$ values. For the lowest cutoff, $\sigma_{\rm F}(\beta)$ slightly exceeds $1\%$, whereas for the two higher cutoffs, it reduces to $0.71\%$ and $0.66\%$, respectively. The reduction in error bar with increasing cutoff is anticipated, as a larger $\Lambda$ permits the inclusion of more modes in both the forward model (“reconstruction”) and the likelihood (BAO inference), thereby providing additional information. We can also compare the behavior of $\sigma_{\rm F}(\beta)$ as a function of $\Lambda$ with the ideal expectation from mode counting, $\sigma_{\rm F}(\beta)\propto N_{\text{mode}}^{-1/2}(\Lambda)$, finding rough agreement. More precisely, the error bar $\sigma_{\rm F}(\beta)$ improves more rapidly than the mode scaling between $\Lambda=0.15\,h\,{\rm Mpc}^{-1}$ and $\Lambda=0.18\,h\,{\rm Mpc}^{-1}$, while between $\Lambda=0.18\,h\,{\rm Mpc}^{-1}$ and $\Lambda=0.2\,h\,{\rm Mpc}^{-1}$, it shrinks at a lower rate than estimated from mode counting. Note however that the simple mode counting comparison ignores the fact that $\partial P_{\rm L}(k,\beta)/\partial\beta$ also depends strongly on $k$ in an oscillatory way.
@@
Figure 3: Field-level BAO inference results for Mock A. The left panel shows the inferred BAO scale relative to ground truth obtained using Lagranigan bias for sampling. FreeIC are represented using circle marker and FixedIC using a square. On the right we show the 68% CL error bar on the BAO scale, $\sigma_{\rm F}(\beta)$, as a function of cutoff $\Lambda$.
@@
Figure 4: Field-level BAO inference results for Mock B. The left panel displays the inferred BAO scale relative to the ground truth, where the BAO scale is sampled alongside the initial conditions. The Lagrangian bias points, computed for the same values of $\Lambda$, have been slightly displaced horizontally for better visibility. On the right, we show the 68% CL error bar, $\sigma_{\rm F}(\beta)$, as a function of the cutoff $\Lambda$. The results obtained using the Lagrangian bias are depicted in blue, while those obtained using Eulerian bias are shown in orange.
@@
Turning our attention to Mock B, which is generated using Eulerian bias, we performed inferences with both Eulerian and Lagrangian bias. The left panel of Fig. 4 shows the inferred $\beta$ for both Eulerian and Lagrangian bias cases. In the instance of Eulerian bias, we observe that the residual systematic bias remains below 1% across all considered cutoffs. Additionally, the ground truth value of $\beta$ is consistently recovered within $1\sigma$. Conversely, when using the Lagrangian bias model for inference, a slightly higher systematic bias is evident. Specifically, for $\Lambda=0.15\,h\,{\rm Mpc}^{-1}$ and $0.18\,h\,{\rm Mpc}^{-1}$, this bias is around 1%, while for $\Lambda=0.2\,h\,{\rm Mpc}^{-1}$, it increases to 1.7%. This is consistent with what we found for Mock A, and would likely be solved by going to a higher order in bias in the inference.
@@
Figure 5: Parameter posteriors for the FreeIC inference in the case of Mock B. Left panel (a) represents the Eulerian bias model while the right panel (b) represents the Lagrangian bias. The dotted gray line indicates the ground truth value $\beta_{0}$ in each case. The inference was performed at $\Lambda=0.18\,h\,{\rm Mpc}^{-1}$ in both cases.
In the right panel of Fig. 4, we show the inferred 68% CL error bar $\sigma_{\rm F}(\beta)$ for Mock B. Comparing the inferred error bars, $\sigma_{\rm F}(\beta)$, between the Eulerian and Lagrangian models shows that the error bar is smaller in the case of the Lagrangian bias model. This difference is likely attributable to the construction of bias operators in these models. In the Eulerian bias model, bias operators are constructed using the filtered matter field, $\delta_{\Lambda}$. This second filter (in addition to the one applied to the initial conditions) removes some mode-coupling contributions that are under control and kept in the Lagrangian bias model, which only filters the initial conditions (both bias models of course use the same likelihood filter). The downside of the Lagrangian approach is the higher computational cost due to the additional density assignments needed.
In Fig. 5, we show the parameter posteriors for the Mock B at the cutoff $\Lambda=0.18\,h\,{\rm Mpc}^{-1}$. The two panels correspond to the Eulerian and Lagrangian biases, respectively. Mock B was generated at $\Lambda_{0}=0.3\,h\,{\rm Mpc}^{-1}$ with $\sigma_{\epsilon}=0.9$. From Eq. (2.19), we find that this corresponds to $\sigma_{\epsilon}\approx 0.41$ at $\Lambda=0.18\,h\,{\rm Mpc}^{-1}$. We see that the inferred value of $\sigma_{\epsilon}$ is indeed consistent with the expected value within errors. However, we notice that $\sigma_{\epsilon,2}$ is now larger than zero, so that the effective noise level at $k\simeq\Lambda$ is increased. This elevation in noise level is due to the fact that the inference is performed at $\Lambda<\Lambda_{0}$. Since the model is nonlinear, the modes between $\Lambda$ and $\Lambda_{0}$ which are integrated out lead to additional effective noise contributions, as shown explicitly in [24] and at the level of renormalization-group equations in [45].
6 Comparing field-level inference to the BAO reconstruction approach